OpenStudy (anonymous):

What is the slope of a line that passes through the point (−1, 1) and is parallel to a line that passes through (3, 6) and (1, −2)?

OpenStudy (anonymous):

i think you have to find the slope first

OpenStudy (anonymous):

For lines to be parallel, they have to have the same slope. So if you find the slope of the second line, you will find the slope of the first line. The equation to do this is: m = y1 - y2 / x1 - x2 So plug in the numbers from your question. m = 6 - (-2) / 3 - 1 = 8 / 2 = 4 The slope of the second line is 4, therefore the slope of the first line is also 4. If you had to find the slopes of both lines to compare them, you would use the equation I showed earlier twice, once for each line.

OpenStudy (anonymous):

Also, I should mention that in the equation: m = y1 - y2 / x1 - x2 y1 can be either the first y value or the second, it doesn't matter. BUT you have to use the same one(either first or second) for the x value that you used for the y value or the equation won't work. ^Sorry if that was confusing, kind of hard to explain.